Circular No. 8507

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bf IN MCG +00-27-5 L. A. G. Monard reports his discovery of an apparent supernova (mag

∼ 18.0 ± 0.3) on unfiltered CCD images taken on Apr. 5.722 and 6.715 UT with the 0.30-m reflector at the Bronberg Observatory near Pretoria, South Africa. The new object is located at α = 10^{h}23^{m}56^{s}.99, δ = −3^{o}11′29′′.3 (equinox 2000.0), which is 13′′ east and 32′′ south of the nucleus of MCG +00-27-5. Nothing is visible at this position on the Digitized Sky Survey (limiting red mag 20.5) or on an image taken by Monard on 2005 Mar. 12.80 (limiting red mag 18.5). Further to IAUC 8504, M. Moore and W. Li independently report the LOSS discovery of SN 2005bf (at mag 17.8) on an unfiltered KAIT image taken on Apr. 6.3), providing position end figures 57^{s}.27, 28′′.6 (and offset 11′′.7 east and 32′′.6 south of the nucleus of MCG +00-27-5). The new object was also marginally detected (at mag ∼ 18.8) on an earlier KAIT image taken on Mar. 30.31, while a KAIT image taken on Mar. 15.31 showed nothing at this position (limiting mag 19.5).

SUPERNOVA 2005az IN NGC 4961 Note that the host galaxy for SN 2005az appears to be NGC 4961 (not

4960 as stated on IAUC 8503). Additional unfiltered CCD magnitudes reported by W. Li (LOSS/KAIT) and T. Puckett and D. Moon (cf. IAUC 8504) for SN 2005az: Mar. 16.43 UT, [19.0 (Li); Apr. 4.22, 16.7 (Puckett); 5.39, 16.3 (Li). SN 2005az was also marginally detected on an KAIT earlier image taken on Mar. 27.41 under poor conditions.

SUPERNOVA 2005aw IN IC 4837A N. Morrell, M. Hamuy, G. Folatelli, and M. Roth, Carnegie Supernova

Project, report that a spectrogram (range 380–930 nm) of SN 2005aw (cf. IAUC 8499), obtained on Apr. 3.33 UT with the Las Campanas 2.5-m du Pont Telescope (+ WFCCD spectrograph), reveals this to be a type-Ic supernova, several days past maximum. The spectrum, dominated by strong Ca II 857.9-nm, along with O I 777.3-nm, Si II 635.5-nm, Na I 589.2-nm, Fe II 508.3-nm, and Mg II 448.1-nm, is very similar to those of supernovae 1994I and 1987M at 10–11 days after maximum light. An equivalent width of 0.13 nm was measured for the interstellar Na I D line arising in the host galaxy.

2005 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green